Current Time: 15:26:04 EST
 

Archive for January, 2017

  • Planned maintenance for mysql1001-1015, pgsql1001-1002, pgsqlcp7, ns13, ns14, ss10 – Resolved

    We are planning on a continuation of maintenance on several servers. This maintenance is to correct storage system issues. This will only affect the customer’s ability to connect to the server for 15-20 minutes. We already moved mail1001 through mail1015 last week, and the downtime on average was between 5 and 10 minutes each. Maintenance […]

  • web519 – Down – Resolved

    Sites hosted on the web519 server are currently inaccessible.  Our systems administration team is currently working on correcting the problem.

  • Denial of Service Attack: iis servers on Control Panels 6, 7, 8 – Resolved

    We are experiencing significant packet loss for the networks assigned to the Windows Servers for control panels 6, 7, and 8; this is due to a Denial-of-Service attack.  Our administration department has been notified and are working on the issue.  We will post updates and information as we learn more.

  • Control Panels 6, 7, 8 – Network Down – Resolved

    We are experiencing significant packet loss for the networks assigned to the servers for control panels 6, 7, and 8.  Our administration department has been notified and are working on the issue.  We will post updates and information as we learn more.

  • web145 Web Server Maintenance – Completed

    Web145 is currently experiencing an issue with the backup drive.  We will need to take the server down in order to exchange the hardware.  This maintenance will begin at 12:40PM EST.

  • web414 Slow – Resolved

    web414 is experiencing a currently-unknown problem causing sites on that server to run slowly.

  • CP10 Mail Server Moves – Resolved

    Beginning tonight we will be moving mail1001 through mail1015 to the new pool as mentioned in Outage  Post Mortem status post. We will begin the migration starting with mail1015 and working backwards until all of the mentioned servers have been moved.

  • iis311 Maintenance – Resolved

    Server was rebooted to check and fix the consistency of the file system.  It was determined the logical file structure is corrupted.  Running file system check to resolve. We have had continued problems with the RAID controller. As a result, it has corrupted the file system and the server crashes while running the file system […]

  • 10-14-10 Outage Post Mortem

    Post-Mortem Thursday, October 14th 11:00pm [all times EST]:  Planned maintenance initiated to install two new switches, one into each redundant stack, in our CP9&CP10 environment in preparation for a new enclosure.  Procedure already implemented with no customer impact in similar environments, two-hour maintenance window requested in case of necessary reboot, but not anticipated to be […]

  • iis532 maintenance – Complete

    Asp.net updates installation. Asp.net updates installation is needed in order to resolve recent issues with customers’ asp.net applications.

  • mysql1005, mysql1007, and mysql1011 – ongoing repair – Resolved

    MySql 1005, 1007, & 1011 are still undergoing repairs from work started before the restart of the SAN storage.  These servers are again being worked on, ETA 3-4 hours for 1005 & 1007, ETA pending for 1011.

  • CP9/CP10 mail and database servers down/CP7 Management affected too – Resolved

    At this time web servers on CP9 and CP10 are up, but mail and database servers are unable to connect to the SAN storage, which is causing them to be unreachable.  Our sysadmins have determined that the problem is with the amount of connections to the SAN (storage) being exhausted, most likely due to the […]

  • Mail408 migrating to new hardware – Complete

    Mail408 is being migrated to new hardware. All services on this server will be down while content is copied from the old mail408 hardware to the new hardware. Service should return within 2 hours. No mail will be lost while the server is down, but it will be delayed. Update 12:56AM: The migration is complete […]

  • CP9 Remaining Servers Down – Resolved

    There are a few  servers still experiencing issues.  If your services are down and your account is not located on one of these servers, contact support as soon as possible: mail905 – Server is online. web919 – Server is online. web936 – Server is online. web938 – Server is online. web947 -Server is online. We […]

  • CP10 Remaining Servers Down – Resolved

    In order to restore full service to CP10, all servers as well as cp services on CP10 must be restarted.  This process has caused all servers on CP10 to be down.  Servers on CP10 will be coming online as they boot over the next two hours. The only CP10 servers that are still down are: […]

 
© 2011 IX Web Hosting.